OpenStudy (anonymous):

PLEASE HELP!! Solve for x: 4(2x-1) = 2x +35 1:3.9 2:6.5 3:6 4:39

jimthompson5910 (jim_thompson5910):

4(2x-1) = 2x +35 8x-4 = 2x +35... distribute 8x - 4 - 2x = 35 ... Subtract 2x from both sides. 8x - 2x = 35 + 4 ... Add 4 to both sides. I'll let you finish

OpenStudy (anonymous):

So do i add four to -2x or 8x? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

8x - 2x simplifies to what?

OpenStudy (anonymous):

6x @jim_thompson5910

jimthompson5910 (jim_thompson5910):

35 + 4 simplifies to _____

OpenStudy (anonymous):

39 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so we now have 6x = 39 therefore, x = ???

OpenStudy (anonymous):

6.5 ;) lol thank you so much ;) it helped alot @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thanks ;) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome